OpenStudy (anonymous):

Can I please get some help :/ Prove the following identity, showing each step in your work. sin (A + B) – sin (A – B) = 2 cos A sin B

terenzreignz (terenzreignz):

You know the identity for the sum of two angles? sin(x + y) = sin x cos y + cos x sin y

OpenStudy (anonymous):

Oh, yea that looks familiar.

OpenStudy (hba):

from the double formula: sin (A + B) = sin A cos B + sin B cos A and sin (A - B) = sin A cos B - sin B cos A if we are going to substitute these to the given equations it will give us Sin (A+B) - Sin (A-B) = [sin A cos B + sin B cos A] - [sin A cos B - sin B cos A] simplify: Sin (A+B) - Sin (A-B) = sin A cos B - sin A cos B + sin B cos A + sin B cos A Sin (A+B) - Sin (A-B) = 2 sin B cos A

OpenStudy (anonymous):

Dang. Thanks hba :)

OpenStudy (hba):

ur welcome @IloveCharlie